### v1

# Claim Verification in the Age of Large Language Models: A Survey

Abstract
The large and ever-increasing amount of data available on the Internet coupled with the laborious task of manual claim and fact verification has sparked the interest in the development of automated claim verification systems. [^1] We use the terms claim verification and fact verification interchangeably given the overlap between the two concepts. Several deep learning and transformer-based models have been proposed for this task over the years. With the introduction of Large Language Models (LLMs) and their superior performance in several NLP tasks, we have seen a surge of LLM-based approaches to claim verification along with the use of novel methods such as Retrieval Augmented Generation (RAG). In this survey, we present a comprehensive account of recent claim verification frameworks using LLMs. We describe the different components of the claim verification pipeline used in these frameworks in detail including common approaches to retrieval, prompting, and fine-tuning. Finally, we describe publicly available English datasets created for this task.
Introduction
Figure 1: Comparison of claim verification systems between NLP-based (traditional) and LLM-based for claim veracity.
Figure 2: The archetypal claim verification pipeline and its main components.
False information is widely present on social media and on the Web motivating the development of automated fact verification systems (Guo, Schlichtkrull, and Vlachos 2022). The introduction of LLMs has provided malicious actors with sophisticated ways of creating and disseminating false information. Recent election cycles saw a large number of claims spread across both social media and news platforms alike (Dmonte et al. 2024). Similarly, during the COVID-19 pandemic, many claims were spread across social media platforms. Many of these claims were factually inaccurate which led to the spread of misinformation (Zhou et al. 2023).
Fact-checking potentially false claims is an essential moderation task to reduce the spread of misinformation. Several organizations such as FactCheck, PolitiFact, NewsGuard, and Full Fact perform manual fact-checking to verify claims in different domains. However, this has been regarded as a laborious task that requires domain expertise (Adair et al. 2017; Hanselowski 2020) and it has become less and less feasible due to the sheer volume of misinformation that can be generated by humans and AI models. Automated fact-checking has become an increasingly popular approach to verify the veracity of claims in a given text. [^2] We acknowledge that claim verification models can also be applied to other modalities of data (e.g., images). In this survey, we address models applied to text only. There are several steps involved in the fact-verification pipelines, but the three main components are claim detection, evidence retrieval, and veracity prediction. Models used in these steps have followed the general methodological developments of the field and we have thus observed an increase in the use of LLMs for claim verification (Zhang and Gao 2023; Wang and Shu 2023; Quelle and Bovet 2024). Figure 1 shows an example claim veracity scenario and compares the architectural differences between NLP-based (traditional) and LLM-based claim verification systems. Compared to traditional systems which use NLP-based models, fact verification are less prone to error propagation, and provide explainable/interpretable generations/justification when verifying claims.
Despite this, LLMs are pre-trained on very large collections of texts and are prone to hallucinations, often generating texts containing incorrect information. Furthermore, such models can be used to generate misinformation at scale (Chen and Shu 2023; Zhou et al. 2023; Dmonte et al. 2024) and therefore they can be exploited by malicious actors to generate and spread wrong and factually incorrect information at an unprecedented rate (Pan et al. 2023c). Furthermore, using these pre-trained models for fact-verification may generate incorrect veracity labels, as the models may also rely on obsolete information to assess the veracity of a claim. Approaches like RAG (Gao et al. 2023), which aid the models in their decision-making abilities are used in this task for the model to get access to the most recent information during the fact-verification task.
A few general automated claim verification surveys have been published over the years (Zeng, Abumansour, and Zubiaga 2021; Bekoulis, Papagiannopoulou, and Deligiannis 2021; Guo, Schlichtkrull, and Vlachos 2022) as well as a couple of surveys focusing on particular aspects of the task such as explainability (Vallayil et al. 2023) and applications to specialized domains such as scientific texts (Vladika and Matthes 2023). However, all past related surveys, lack consideration for LLM-based approaches or focus on specific sub-tasks of the pipeline (Panchendrarajan and Zubiaga 2024). In this paper, we fill this important gap by surveying LLM-based frameworks proposed in recent years. To the best of our knowledge, this is the first survey to explore claim verification with LLMs. We expect it to be a valuable resource to researchers in the field opening exciting new avenues for future research.
Search Criteria
We search various well-known repositories [^3] ACM: portal.acm.org. IEEE Xplore: ieeexplore.ieee.org. Scopus: scopus.com. ACL: aclanthology.org. Web of Science: isiknowledge.com. Springer: link.springer.com. ArXiv: arxiv.org. CEUR: https://ceur-ws.org/. of scientific articles to collect the papers that serve as primary sources for this survey. We apply queries including LLMs, claim verification, fact verification and related keywords on these repositories. We focus primarily on the ACL Anthology, the ACM Digital Library, and IEEE Xplore, and proceedings of related conferences such as AAAI and IJCAI. We further search on Scopus, Springer Link, and Science@Direct, and ArXiv.
We collected and reviewed over 100 papers based on our search terms and filter papers based on the following search criteria: (i) since the scope of our review is on text-based LLM claim verification systems, we omit papers related to multimodal approaches (e.g., text-to-images), our papers from other modalities in claim verification (e.g., images, graph-based systems). (ii) since our review focuses on LLM-based systems for claim veracity, we omit papers related to claim identification, claim detection, and/or detecting LLM-generated content. Based on our filtering criteria, we collected a total of 49 papers related to LLM-based approaches for veracity labeling. The papers on topic LLM-based veracity labeling have been published primarily in the ACL Anthology and ACM Digital Library, but also in other repositories such as the IEEE Xplore. We release this curated list to be publicly available.
Figure 3: Claim verification pipeline using large language models. Instead of traditional automatic claim verification, this pipeline creates a prompt from the retrieved evidence and the input claim as input to the LLM to generate a label, sentence evidence, and/or explanation of its response.
Claim Verification Pipeline
Figure 2 shows an archetypal claim verification pipeline consisting of the following modules: claim detection, claim matching, claim check-worthiness, document/evidence retrieval, rationale/sentence selection, veracity label prediction, and explanation/justification generation. Many proposed systems make use of only some of these modules. The following subsections describe each of these modules in detail. Finally, as described in Panchendrarajan and Zubiaga (2024), performance evaluation in these sub-tasks is usually carried out using well-established automatic evaluation metrics used in text classification such as Precision, Recall, and F-score.
Claim Detection
Input texts may contain one or more statements but not all statements are claims. Given the input text, a claim detection module is designed to identify all the statements containing a claim. For example, the statement ’I loved the movie Oppenheimer.’ is an opinion, whereas the statement ’The COVID-19 pandemic started in Texas.’ contains a claim.
Check-worthy Claim Identification
Not all the identified claims are check-worthy. In the check-worthy claim identification sub-task, given an input claim, the task is to identify the claims that include real-world assertions and that may need to be verified (Hassan, Li, and Tremayne 2015; Nakov et al. 2021). This is a subjective task and it relies on the factors like popularity of the claim, public interest in determining the veracity of the claim, etc. For example, the claim ’The President met the State Governor to discuss the infrastructure deal.’ is less check-worthy than the claim ’Drinking salt water cures COVID.’.
Claim Matching
The identified check-worthy claims can be matched to the previously fact-checked claims. Given an input claim and a database of previously fact-checked claims, claim matching is used to determine if the input claim is previously fact-checked and exists in the database (Shaar et al. 2020; Nakov et al. 2021). This can help avoid the next steps in the pipeline and a veracity label can be predicted directly.
Document/Evidence Retrieval
If the input claim does not exist in the database of fact-checked claims, the claim needs to be verified. In the document or evidence retrieval sub-task, all the relevant documents related to the input claim are extracted, either from an external database or through an internet search/information retrieval(Chen et al. 2017). A threshold of how many documents are to be retrieved can be predetermined.
Rationale/Sentence Selection
All the information in the retrieved documents is not relevant to the claim. Hence in the rationale selection task, only the information or evidence most relevant to the claim are selected to be used to predict the veracity label.
Veracity Label Prediction
Once the rationale is selected, the claim along with the rationale and additional features if any, are given as input to a machine learning model. The task of the classifier is to predict a veracity label from among the following three labels; ’SUPPORTED’, ’REFUTED’, or ’NOT ENOUGH EVIDENCE’. In some datasets, the labels can also be ’TRUE’, or ’FALSE’, depending on the specific task.
Explanation/Justification Generation
Recent works have focused on generating explanations for the veracity labels prediction. This specific task is focused on generating natural language justifications or explanations for the prediction considering the claim and evidence to generate these explanations.
LLM Approaches
In this section, we describe the recent advancements that enable LLMs to be robust in fact verification scenarios. Figure 3 shows an example pipeline that encapsulates multiple component modules (i.e., Evidence Retrieval, Prompt Creation, Transfer Learning, and LLM Generation) for verifying claims. Different from traditional fact verification pipelines that select evidence set for verifying claims, LLM-based claim verification conditions generated text based on the concatenated input claim and retrieved evidence. The ability of large, pre-trained LMs augmented with retrieval enables them to perform well on knowledge-intensive tasks such as text generation.
Evidence Retrieval Strategies
RAG models, which have been developed to combat the issue of hallucination in LLMs in knowledge-intensive tasks, have shown success in the scope of fact verification (Lewis et al. 2020; Izacard et al. 2023; Gao et al. 2023; Guan et al. 2023). Early work such as in Lewis et al. (2020) developed a framework that incorporates a retriever to retrieve evidence from an external database such as Wikipedia for conditionally generating veracity labels in fact verification. Authors in Izacard et al. (2023) demonstrate that RAGs perform well on the FEVER shared task (Thorne et al. 2018b)) in few-shot settings, showing approximately 5% improvement over large-scale LLMs such as Gopher (Rae et al. 2021) with significantly fewer parameters (i.e., 11B compared to 280B). Other works consider the optimization of either document ranking (Glass et al. 2022; Chen et al. 2022c; Hofstätter et al. 2023) or input claims (i.e., queries) (Hang, Yu, and Tan 2024) as a crucial step for improving evidence retrieval for veracity labeling. Authors in Hofstätter et al. (2023) use an autoregressive re-ranker to get the most relevant passages from the retriever. These are then passed to the generation model to generate the veracity label.
Despite this, RAG models can often fail when encountering long or complex input claims, causing the model to incorrectly generate veracity labels or evidence sentences. Recent works have addressed this issue by segmenting long claims into smaller sub-claims and performing multiple rounds of retrieval (Khattab, Potts, and Zaharia 2021; Shao et al. 2023). Authors in Khattab, Potts, and Zaharia (2021) present a pipeline for multi-hop claim verification that uses an iterative retriever and neural methods for effective document retrieval and re-ranking.  Shao et al. (2023) show that using a re-ranker to distill knowledge to a retriever helps close the semantic gaps between a query and document passage when verifying claims using iterative RAGs. Other works address the issue of complex claims by using fine-grained retrieval techniques based on claim decomposition (Chen et al. 2023a; Zhang and Gao 2023; Pan et al. 2023b).Chen et al. (2023a) generate sub-questions based on a claim, which a document retriever uses to retrieve relevant documents. A fine-grained retriever retrieves top-k text spans as evidence based on a k-word window and BM25. Authors in Zhang and Gao (2023) decompose a claim into sub-claims and generate questions to verify the sub-claims. External knowledge sources are used to retrieve relevant information to verify the sub-claims and generate a final veracity label. Pan et al. (2023b) follow a programming paradigm, where the claim is broken down into subtasks and the final label is the aggregation of the execution of each subtask. The fact-verification subtask uses external knowledge to retrieve relevant evidence for a claim.
Hang, Yu, and Tan (2024) retrieve evidence based on generated knowledge graphs of evidences. They generate a knowledge graph of user query or input claim and compare it to the database of knowledge graphs to retrieve the most relevant information for claim verification. Authors in Hu et al. (2023) propose a latent variable model that allows the retrieval of the most relevant evidence sentences from a document while removing the irrelevant sentences. This approach reduces the noisy data during the verification process. Stochastic-RAG, an approach proposed by Zamani and Bendersky (2024) uses a stochastic sampling without replacement process for evidence retrieval and selection. This approach overcomes the ranking and selection of the evidence hence optimizing the RAG model. Authors in Zhang et al. (2023) optimize the evidence retrieval process by using feedback from the claim verifier. The divergence between the evidence from a retrieved evidence set provided to the verifier and the gold standard evidence, acts as a feedback signal used to train the retriever. Xu et al. (2024b) propose Search-in-the-Chain, an approach where an LLM generates a reasoning chain and based on the answer to each node in the chain, retrieval can be used to correct an answer or provide additional knowledge. The approach improves the generation accuracy of the LLM.
Prompt Creation Strategies
Text prompting has been shown to be an effective technique for improving the desired output of large-scale generative models. In the context of claim verification, several works investigate prompting strategies, using both manual and automated techniques, for building more robust claim verification systems Zhang and Gao (2023); Li et al. (2023c); Zeng and Gao (2023); Chen et al. (2023b). Authors in (Zhang and Gao 2023) develop a hierarchical prompting technique that enables LLMs to verify multiple sub-claims using a step-by-step approach. The work in Li et al. (2023c) demonstrates a self-sufficient claim verification through prompting instructions on multiple language models.
ProToCo (Zeng and Gao 2023) demonstrates improved claim verification performance of LLMs by leveraging a consistency mechanism to construct variants of the original claim-evidence pair prompt based on three logical relations (i.e., confirmation, negation, uncertainty). Authors in Chen et al. (2023b) develop a unified retrieval framework that employs discrete, continuous, and hybrid prompt strategies for adjusting to various knowledge-intensive tasks such as claim verification. Other works such as the FactualityPrompts (Lee et al. 2022) framework test the output generations of LLMs given an input prompt and use an external database such as Wikipedia to calculate factuality and quality measures compared to the ground truth. Other works aim to improve the LLM’s reasoning abilities by appending the claims with evidence during prompting in the context of claim verification and text generation (Parvez 2024; Dougrez-Lewis et al. 2024).
Transfer Learning Strategies
Fine-Tuning.
Although recent studies show the success of pre-trained LLMs on zero- or few-shot tasks, they often fail to verify real-world claims given their limited internal knowledge. The success of fine-tuning has motivated recent work on claim verification (Chen et al. 2022b; Pan et al. 2021; Zeng and Zubiaga 2024). The work in Chen et al. (2022b) leverages a language model to fine-tune over an external corpus for retrieving passage titles and evidence sentences using constrained beam search. The results showed improved performance against the traditional fact verification pipeline over the FEVER dataset. Other work demonstrates that using GPT models to generate synthetic training data improves the performance of LLMs on various tasks such as fact checking (Tang, Laban, and Durrett 2024) and claim matching (Choi and Ferrara 2024).
Authors in Pan et al. (2021) develop a pipeline for creating a fact verification dataset and fine-tuning a language model by leveraging passages from Wikipedia to generate QA pairs related to claim veracity. The authors show that it can improve state-of-the-art language models in zero-shot settings. The work in (Zeng and Zubiaga 2024) shows that using unlabelled pairwise data to increase the alignment between claim-evidence pairs shows significant improvement of LLM performance in few-shot claim verification tasks. In addition, other recent works leverage techniques such as reinforcement learning to fine-tune models for improving the veracity of claims and supporting evidence Zhang and Gao (2024); Huang et al. (2024). A document-level and question-level retrieval policy is proposed by Zhang and Gao (2024), where the top-k and top-1 documents for the document and question-level policy respectively are used as input to a scoring function for label prediction during training. This approach outperforms retrieval and prompting approaches. Chiang et al. (2024) fine-tune LLMs for their multi-stage fact verification approach. They fine-tune a model to generate answers based on claim-evidence pairs and a set of questions, whereas another model is fine-tuned to verify the claim based on the claim-evidence and question-answer pairs. Authors in Zhu et al. (2023) fine-tune a generation model to generate counterfactuals to train the fact verification model’s performance on out-of-domain claims.
In-Context Learning.
The recent success in the performance of pre-trained LLMs in zero- and few-shot settings is largely attributed to its in-context learning (ICL) abilities (Kojima et al. 2022; Brown et al. 2020). In the scope of claim verification, popular ICL techniques with strong zero- and few-shot performance include chain-of-thought (CoT) reasoning (Wei et al. 2022). The work in Zhao et al. (2024) develops a multi-stage verification pipeline for claim verification based on claim decomposition and self-reflection. An LLM-based verifier module is created using instruction prompting to generate a reasoning analysis among all sub-claims created by the decomposer module. The results suggest that using zero-shot prompting techniques provides better performance in multi-hop claim verification domains such as HOVER and FEVEROUS compared to few-shot prompting and fine-tuning methods. Authors in Kanaani (2024) enable LLMs to generate reasons over retrieved evidence in claim verification using few-shot ICL and the STaR CoT technique inspired by the work of Zelikman et al. (2022). Similar work has leveraged CoT techniques for the purposes of effectively verifying complex claims using reasoning steps (Yao et al. 2023; Ni et al. 2024). On the other hand, HiSS (Zhang and Gao 2023) demonstrates that using prompting techniques for few-shot learning and claim decomposition can substantially improve the performance of CoT models for complex news claim verification. Li et al. (2023b) leverage the ICL capability of LLMs to perform multiple tasks simultaneously. Their approach outperforms or achieves comparable task performance in a zero-shot setting on claim verification datasets.
LLM Generation Strategies
Label and Evidence Generation.
While the majority of claim verification systems predict veracity labels based on the concatenation of the input claim and the set of evidence sentences (Pradeep et al. 2021), recent work has proposed alternate strategies for determining veracity labels as well as selecting/generating evidence pieces. The work in Cao et al. (2024) develops, SERIf, a claim verification pipeline that features an inference module that predicts the veracity label of scientific news articles based on a two-step summarization (i.e., ‘Extractive - Abstractive‘) and evidence retrieval technique. Each summary-evidence pair is fed into the LLM and it produces a binary label, indicating whether the news article is reliable (supported) or unreliable (refuted). Authors in Wadden et al. (2022b) leverage the Longformer transformer model (Beltagy, Peters, and Cohan 2020) that uses a shared encoding over the claim and document abstracts for rationale identification and claim label prediction.
The work in Li et al. (2024a) develops an algorithm that selects the minimal evidence group (MEG) within a set of retrieved candidate documents. The algorithm aims to minimize the redundancy while also selecting the most relevant piece of evidence to prompt the language model. Authors in Chen et al. (2022b) create an LLM claim verification framework in which the authors use the BART model to encode all candidate sentences from the most relevant retrieved documents. The BART decoder serves as an evidence decoder to predict the $g$-th evidence sentence via generation conditioned on the top $k$ retrieved documents and the input claim. Lee et al. (2022) developed a variant of nucleus sampling called, factual-nucleus sampling, in which the top-$p$ sampling pool is selected as a set of sub-words whose cumulative probability exceeds $p$. The authors show that factual-nucleus sampling can improve evidence and label generation without claim verification datasets such as FEVER. Kao and Yen (2024) propose a multi-stage approach, where the evidence sentences are retrieved from articles related to a claim, and arguments are generated by aggregating and reconstructing the evidence. The arguments are refined and passed to a LLM to generate a verification label. Other approaches leverage LLMs reasoning capabilities to generate veracity labels and factual evidence (Cheng, Tan, and Lu 2024; Jafari and Allan 2024; Li et al. 2024b; Fang et al. 2024; Pan et al. 2023a).
Dataset # Instances Source # Labels Reference
AVeriTeC 4,568 Google FactCheck Claim Search API 4 Schlichtkrull, Guo, and Vlachos (2024)
BINGCHECK 3,840 Microsoft Copilot 4 Li et al. (2023c)
Check-COVID 1,504 News websites, CORD-19 3 Wang et al. (2023a)
CLAIMDECOMP 1,000 PolitiFact 6 Chen et al. (2022a)
CLIMATE-FEVER 1,535 Google Search, Wikipedia 3 Diggelmann et al. (2020)
CoVERT 212 Manually created 3 Liu et al. (2024)
COVID-Fact 4,086 Reddit 2 Saakyan, Chakrabarty, and Muresan (2021)
EX-FEVER 60,000 Wikipedia 3 Ma et al. (2024)
ExpertQA 2177 LLM, Google Search 5 Malaviya et al. (2024)
FACTIFY-5WQA 391,041 Public Fact-Verification Datasets 3 Rani et al. (2023)
FactCheck-Bench 678 LLM 4 Wang et al. (2023b)
FactCheckQA 20,871 PolitiFact 3 Bashlovkina et al. (2023)
FACTUALITYPROMPTS 16,000 Wikipedia, LLM 2 Lee et al. (2022)
FAVIQ 188,000 Wikipedia, Existing Datasets 2 Park et al. (2022)
FEVER 185,445 Wikipedia 3 Thorne et al. (2018a)
FEVEROUS 87,026 Wikipedia 3 Aly et al. (2021)
FoolMeTwice 12,968 Wikipedia 2 Eisenschlos et al. (2021)
HealthVer 14,330 Bing Web Search API 3 Sarrouti et al. (2021)
HOVER 26,171 Wikipedia 2 Jiang et al. (2020)
LIAR 12,500 PolitiFact 6 Wang (2017)
MultiFC 34,918 Fact-checking Websites 2-40 Augenstein et al. (2019)
PolitiFact 21,152 PolitiFact 6 Misra (2022)
QuanTemp 15,514 Google Fact Check Tool API 3 Venktesh et al. (2024)
RAWFC 2,012 Snopes 6 Yang et al. (2022)
SciFact 1,400 S2ORC 2 Wadden et al. (2020)
SciFact-Open 279 S2ORC 3 Wadden et al. (2022a)
SciNews 2,400 Wiki Fake News, LLM, Existing Datasets 2 Cao et al. (2024)
SCITAB 1,225 SciGen Dataset (arXiv) 3 Lu et al. (2023)
VitaminC 400,000 Wikipedia 3 Schuster, Fisch, and Barzilay (2021)
WiCE 1,967 Wikipedia 3 Kamoi et al. (2023)
XClaimCheck 16,177 PolitiFact 5 Kao and Yen (2024)
Table 1: Available datasets along with the number of instances, data sources, and references.
Explainable Generation.
Recent studies investigate explainable approaches to improve LLM-based claim verification systems (Wang and Shu 2023; Dammu et al. 2024; Si et al. 2023). Authors in Wang and Shu (2023) present an explainable claim verification framework named FOLK, that leverages the explanation capabilities of LLMs when verifying a claim and justifies the prediction through a summary of its decision process. The work in Dammu et al. (2024) proposes a knowledge graph (KG)-based approach for text verification and evidence attribution. An objective function is used to fine-tune LLMs on evidence attribution based on the input text and retrieved triplets from KG, inducing explanations on claim predictions. While explainable techniques can aid humans in fact-checking, LLMs are prone to incorrect explanations due to hallucinations, causing them to be unreliable in certain claim verification scenarios (Si et al. 2023). Ma et al. (2024) prompt an LLM in a few-shot setting to generate a concise summary of the gathered evidence documents and the input claim. This summary serves as an explanation for the verified claim. Other works in claim verification leverage reasoning techniques such as chain-of-thought (CoT) for enabling the LLM to be interpretable in its decision-making process when verifying claims (Yao et al. 2023; Pan et al. 2023a; Zhao et al. 2024; Kanaani 2024; Ni et al. 2024; Quelle and Bovet 2024; Fang et al. 2024). The authors in Pan et al. (2023a) and Fang et al. (2024) leverage the reasoning ability of the LLMs to generate explanations by using question-guided reasoning and minimizing the inherent model biases respectively.
Evaluation and Benchmarking
Metrics
The F1 score is the most commonly used metric to measure the performance of automatic claim verification systems. Other metrics like Precision, Recall, and Accuracy are also used to evaluate the system performance. Katranidis and Barany (2024) used the error rate between the human and automated fact-verification system to measure the verification accuracy. However, these metrics consider a single pipeline component to evaluate the system’s overall performance. Hence, Thorne et al. (2018a) introduced the FEVER score, a metric that uses both the verification accuracy as well as the evidence retrieval accuracy to compute overall system performance. While these metrics are valuable for assessing the performance of the classification tasks, they are inadequate for evaluating the performance of the non-classification components of the pipeline. Hence, metrics like Recall@k are used to measure the performance of the retrieval task (Pan et al. 2023b; Pradeep et al. 2021), while BLEU and METEOR are used to evaluate the quality of explanations or generated questions and answers. Schlichtkrull, Guo, and Vlachos (2024) propose a new evaluation metric AVeriTeC score that uses METEOR and accuracy metrics, for question-answer-based veracity label prediction systems. Other metrics like Mean Absolute Error (MAE), Expected Calibration Error (ECE), Area Under ROC Curve (AUC-ROC), and Pearson’s Correlation are also used. Most of these metrics are inadequate to evaluate the factual accuracy of the LLM-generated text. Several metrics like FactScore (Min et al. 2023), SAFE (Wei et al. 2024), and VERISCORE (Song, Kim, and Iyyer 2024) are introduced to evaluate the factual accuracy of the LLM-generated text. While these metrics evaluate the accuracy, other metrics and frameworks to evaluate the factual errors in generated text (Lee et al. 2022; Chern et al. 2023) as well as their alignment (Zha et al. 2023) and entailment (Lee et al. 2022) considering the factuality have been proposed.
Datasets
The fundamental resource for training and evaluating claim verification systems is datasets containing annotated texts. As most research in this area deals with English data, we collect information about all publicly available English datasets used in the papers discussed in this survey and present them in Table 1.
Several general-domain datasets have been released over the years. Different online data sources and websites are used to extract facts and claims including Wikipedia (Thorne et al. 2018a; Jiang et al. 2020; Diggelmann et al. 2020; Eisenschlos et al. 2021; Schuster, Fisch, and Barzilay 2021; Kamoi et al. 2023), due to the extensive amount of information available spanning various topics and domains, and online fact-checking websites like PolitiFact (Wang 2017; Augenstein et al. 2019; Kao and Yen 2024). Factually incorrect claims are shared through social media channels like X, Reddit, etc. Saakyan, Chakrabarty, and Muresan (2021) introduced the COVID-Fact dataset consisting of claims extracted from Reddit posts. Several datasets for scientific fact verification have also been introduced over the years (Wadden et al. 2020, 2022a; Lu et al. 2023). Most fact-verification datasets rely on unstructured textual evidence. Hence, a few datasets with structured data sources as evidence have been introduced (Aly et al. 2021; Lu et al. 2023). While a large number of datasets are focused on only veracity labels, some datasets were deloped to address explainability as well (Chen et al. 2022a; Yang et al. 2022; Ma et al. 2024; Rani et al. 2023; Schlichtkrull, Guo, and Vlachos 2024).
Most of the claim-verification datasets include claims extracted from available information sources. LLMs are widely used to generate content, given their superior text-generation capabilities. These models often generate factually incorrect information. Efforts to identify such non-factual text have been undertaken. However, existing claim-verification datasets can be inadequate for this task due to the linguistic variations between the human- and LLM-generated text. To overcome this issue, LLM-generated claim verification datasets have been introduced (Li et al. 2023c; Cao et al. 2024). While these datasets are used to evaluate the automatic claim verification systems, there is a need to evaluate the factual accuracy of the LLM-generated text. Hence, a few datasets to evaluate the LLM generation’s factual accuracy have been introduced (Lee et al. 2022; Wang et al. 2023b; Malaviya et al. 2024).
Shared Tasks
Shared tasks are competitions where participating teams develop systems to solve a task using a common benchmark dataset. There have been multiple shared tasks on the claim and fact-checking organized over the years like the Fact Extraction and Verification (FEVER) (Thorne et al. 2018b), TabFact [^4] https://competitions.codalab.org/competitions/21611, CLEF 2020 CheckThat! (Barrón-Cedeno et al. 2020), SCIVER (Wadden and Lo 2021), SEM-TAB-FACT (Wang et al. 2021), FACTIFY-5WQA [^5] https://defactify.com/factify3.html, and more recently AVeriTeC [^6] https://fever.ai/task.html. While there have been various techniques like question-answer generation as a precursor task to label prediction as in the AVeriTeC shared task, all these shared tasks are centered on predicting a veracity label for a claim. Given the LLM’s tendency to hallucinate generating plausible yet factually inaccurate text, there is a need to organize shared tasks to evaluate the factual accuracy of LLM-generated content.
Open Challenges
Handling Irrelevant Context
Retrieved evidence may be irrelevant, which is a challenge for LLMs, as they may not be trained to ignore such evidence. The lack of robustness to the noise can cause the LLM to produce misinformation and incorrect verification. Recent research on open-domain question answering shows that external knowledge relevant to the task can aid the model performance, however, irrelevant context can also lead the model to make inaccurate predictions (Petroni et al. 2020; Shi et al. 2023; Li et al. 2023a; Yu et al. 2023). For the fact-verification task, recent works have proposed techniques to identify the most relevant context while aiding the veracity label prediction and thus improving the overall system performance (Wang et al. 2023c; Yoran et al. 2024; Xia et al. 2024). However, more work on the topic is required to verify the effectiveness of this design approach across multiple fact verification domains.
Handling Knowledge Conflicts
Fact-verification approaches reliance on retrieved evidence can cause knowledge conflicts in LLM-based approaches where retrieved evidence as the internal parameters of the pre-trained LLM may conflict with the external knowledge. This causes the LLM to ignore the retrieved evidence and produces hallucinations. Xu et al. (2024a) provides an in-depth analysis of this scenario. Works by (Li et al. 2023a; Neeman et al. 2023; Mallen et al. 2023; Longpre et al. 2021; Chen, Zhang, and Choi 2022) introduce approaches to avoid knowledge conflicts and mitigate hallucinations for question-answering. Expanding this work to fact-verification, especially to the approaches that use LLMs, is vital for an effective verification process.
Multilinguality
Most automated claim verification approaches rely on English datasets. Furthermore, there are limited multilingual fact-verification datasets (Gupta and Srikumar 2021; Kazemi et al. 2022; Pikuliak et al. 2023; Singh et al. 2023). This hinders the development of approaches for multilingual fact-verification, which achieve the best performance when trained on language-specific datasets (Panchendrarajan and Zubiaga 2024).
Conclusion
We presented a survey on LLM approaches to claim verification. To the best of our knowledge, this is the first claim verification survey to focus exclusively on LLM approaches thus filling an important gap in the literature. We have described each of the sub-tasks of the typical claim verification pipeline and discussed various LLM-based approaches used in this task. Finally, we have also described publicly available English datasets providing important information to new and seasoned researchers on this topic.
Advances in LLM development will likely continue improving the quality of claim verification systems. We hope this survey motivates future research on this topic taking advantage of recently-proposed LLMs, RAG methods, etc. Claim verification is a vibrant research topic and we see multiple open research directions as described in the next sub-section.

### v2

@@
 We describe the different components of the claim verification pipeline used in these frameworks in detail including common approaches to retrieval, prompting, and fine-tuning.
-Finally, we describe publicly available English datasets created for this task.
-Introduction
-Figure 1: Comparison of claim verification systems between NLP-based (traditional) and LLM-based for claim veracity.
-Figure 2: The archetypal claim verification pipeline and its main components.
-False information is widely present on social media and on the Web motivating the development of automated fact verification systems (Guo, Schlichtkrull, and Vlachos 2022).
+Finally, we describe publicly available English datasets for this task.
+1 Introduction
+False information is widely present on social media and on the Web motivating the development of automated fact verification systems Guo et al.
+(2022).
 The introduction of LLMs has provided malicious actors with sophisticated ways of creating and disseminating false information.
-Recent election cycles saw a large number of claims spread across both social media and news platforms alike (Dmonte et al.
-2024).
-Similarly, during the COVID-19 pandemic, many claims were spread across social media platforms.
-Many of these claims were factually inaccurate which led to the spread of misinformation (Zhou et al.
-2023).
+Recent election cycles has seen a large number of claims spread across social media and news platforms alike Dmonte et al.
+(2024).
+Similarly, during the COVID-19 pandemic, many factually inaccurate claims were spread Zhou et al.
+(2023).
 Fact-checking potentially false claims is an essential moderation task to reduce the spread of misinformation.
-Several organizations such as FactCheck, PolitiFact, NewsGuard, and Full Fact perform manual fact-checking to verify claims in different domains.
-However, this has been regarded as a laborious task that requires domain expertise (Adair et al.
-2017; Hanselowski 2020) and it has become less and less feasible due to the sheer volume of misinformation that can be generated by humans and AI models.
+Organizations such as FactCheck, PolitiFact, NewsGuard, and Full Fact perform manual fact-checking to verify claims in different domains.
+However, this is a laborious task requiring domain expertise Adair et al.
+(2017); Hanselowski (2020), which has become increasingly infeasible due to the sheer volume of misinformation generated by humans and AI models.
 Automated fact-checking has become an increasingly popular approach to verify the veracity of claims in a given text.
@@
 In this survey, we address models applied to text only.
-There are several steps involved in the fact-verification pipelines, but the three main components are claim detection, evidence retrieval, and veracity prediction.
-Models used in these steps have followed the general methodological developments of the field and we have thus observed an increase in the use of LLMs for claim verification (Zhang and Gao 2023; Wang and Shu 2023; Quelle and Bovet 2024).
-Figure 1 shows an example claim veracity scenario and compares the architectural differences between NLP-based (traditional) and LLM-based claim verification systems.
-Compared to traditional systems which use NLP-based models, fact verification are less prone to error propagation, and provide explainable/interpretable generations/justification when verifying claims.
-Despite this, LLMs are pre-trained on very large collections of texts and are prone to hallucinations, often generating texts containing incorrect information.
-Furthermore, such models can be used to generate misinformation at scale (Chen and Shu 2023; Zhou et al.
-2023; Dmonte et al.
-2024) and therefore they can be exploited by malicious actors to generate and spread wrong and factually incorrect information at an unprecedented rate (Pan et al.
-2023c).
-Furthermore, using these pre-trained models for fact-verification may generate incorrect veracity labels, as the models may also rely on obsolete information to assess the veracity of a claim.
-Approaches like RAG (Gao et al.
-2023), which aid the models in their decision-making abilities are used in this task for the model to get access to the most recent information during the fact-verification task.
-A few general automated claim verification surveys have been published over the years (Zeng, Abumansour, and Zubiaga 2021; Bekoulis, Papagiannopoulou, and Deligiannis 2021; Guo, Schlichtkrull, and Vlachos 2022) as well as a couple of surveys focusing on particular aspects of the task such as explainability (Vallayil et al.
-2023) and applications to specialized domains such as scientific texts (Vladika and Matthes 2023).
-However, all past related surveys, lack consideration for LLM-based approaches or focus on specific sub-tasks of the pipeline (Panchendrarajan and Zubiaga 2024).
+There are several steps involved in the fact-verification pipelines; the three main components are claim detection, evidence retrieval, and veracity prediction.
+Models used in these steps have followed the general methodological developments of the field and we have thus observed an increase in the use of LLMs for claim verification Zhang and Gao (2023); Wang and Shu (2023); Quelle and Bovet (2024).
+Figure 1 shows an example claim veracity scenario and compares the architectural differences between traditional NLP-based vs LLM-based claim verification systems.
+The latter are less prone to error propagation, and provide additional justifications when verifying claims.
+Figure 1: Comparison of claim verification systems between NLP-based (traditional) and LLM-based for claim veracity.
+Despite this, LLMs are pre-trained on very large text collections texts and are prone to hallucinations, often generating texts containing incorrect information.
+Further, such models can be used to generate misinformation at scale Chen and Shu (2023); Zhou et al.
+(2023); Dmonte et al.
+(2024) and can therefore be exploited by malicious actors to spread factually incorrect information at an unprecedented rate Pan et al.
+(2023c).
+Furthermore, using these pre-trained models for fact-verification may generate incorrect veracity labels, as the models may also rely on obsolete information.
+Approaches like RAG Gao et al.
+(2023) are used for the model to retrieve the most recent information during fact verification.
+A few general automated claim verification surveys have been published Zeng et al.
+(2021); Bekoulis et al.
+(2021); Guo et al.
+(2022) as well as a couple of surveys on particular aspects of the task such as explainability Vallayil et al.
+(2023) and applications to specialized domains such as scientific texts Vladika and Matthes (2023).
+However, all past related surveys lack consideration for LLM-based approaches or focus on specific sub-tasks of the pipeline Panchendrarajan and Zubiaga (2024).
 In this paper, we fill this important gap by surveying LLM-based frameworks proposed in recent years.
 To the best of our knowledge, this is the first survey to explore claim verification with LLMs.
-We expect it to be a valuable resource to researchers in the field opening exciting new avenues for future research.
-Search Criteria
-We search various well-known repositories [^3] ACM: portal.acm.org.
+2 Search Criteria
+We search popular repositories [^3] ACM: portal.acm.org.
 IEEE Xplore: ieeexplore.ieee.org.
@@
 CEUR: https://ceur-ws.org/.
-of scientific articles to collect the papers that serve as primary sources for this survey.
-We apply queries including LLMs, claim verification, fact verification and related keywords on these repositories.
-We focus primarily on the ACL Anthology, the ACM Digital Library, and IEEE Xplore, and proceedings of related conferences such as AAAI and IJCAI.
-We further search on Scopus, Springer Link, and Science@Direct, and ArXiv.
-We collected and reviewed over 100 papers based on our search terms and filter papers based on the following search criteria: (i) since the scope of our review is on text-based LLM claim verification systems, we omit papers related to multimodal approaches (e.g., text-to-images), our papers from other modalities in claim verification (e.g., images, graph-based systems).
-(ii) since our review focuses on LLM-based systems for claim veracity, we omit papers related to claim identification, claim detection, and/or detecting LLM-generated content.
-Based on our filtering criteria, we collected a total of 49 papers related to LLM-based approaches for veracity labeling.
+of scientific articles to collect the papers that serve as primary sources for this survey, using the terms LLMs, claim verification, fact verification and related keywords.
+We focus primarily on the ACL Anthology, ACM Digital Library, IEEE Xplore, and proceedings of related conferences such as AAAI and IJCAI.
+We further search on Scopus, Springer Link, Science@Direct, and ArXiv.
+Based on our search terms, we identify over 100 papers, and filter them as follows: (i) since the scope of our review is text-based LLM claim verification systems, we omit papers related to multimodal approaches (e.g., text-to-images), and papers investigating other modalities (e.g., images, graph-based systems); (ii) since we focus on LLM-based systems for claim veracity, we omit papers on claim identification, claim detection, and/or detecting LLM-generated content.
+This results in a total of 49 papers related to LLM-based approaches for veracity labeling.
 The papers on topic LLM-based veracity labeling have been published primarily in the ACL Anthology and ACM Digital Library, but also in other repositories such as the IEEE Xplore.
-We release this curated list to be publicly available.
-Figure 3: Claim verification pipeline using large language models.
-Instead of traditional automatic claim verification, this pipeline creates a prompt from the retrieved evidence and the input claim as input to the LLM to generate a label, sentence evidence, and/or explanation of its response.
-Claim Verification Pipeline
-Figure 2 shows an archetypal claim verification pipeline consisting of the following modules: claim detection, claim matching, claim check-worthiness, document/evidence retrieval, rationale/sentence selection, veracity label prediction, and explanation/justification generation.
-Many proposed systems make use of only some of these modules.
-The following subsections describe each of these modules in detail.
-Finally, as described in Panchendrarajan and Zubiaga (2024), performance evaluation in these sub-tasks is usually carried out using well-established automatic evaluation metrics used in text classification such as Precision, Recall, and F-score.
+3 Claim Verification Pipeline
+Figure 2 shows a typical claim verification pipeline consisting of: claim detection, claim matching, claim check-worthiness, document/evidence retrieval, rationale/sentence selection, veracity label prediction, and explanation/justification generation.
+Many systems only make use of some of these modules.
+Figure 2: A typical claim verification pipeline
 Claim Detection
-Input texts may contain one or more statements but not all statements are claims.
+Input texts may contain one or more statements, not all of which are claims.
 Given the input text, a claim detection module is designed to identify all the statements containing a claim.
@@
 Check-Worthy Claim Identification
-Not all the identified claims are check-worthy.
-In the check-worthy claim identification sub-task, given an input claim, the task is to identify the claims that include real-world assertions and that may need to be verified (Hassan, Li, and Tremayne 2015; Nakov et al.
-2021).
-This is a subjective task and it relies on the factors like popularity of the claim, public interest in determining the veracity of the claim, etc.
-For example, the claim ’The President met the State Governor to discuss the infrastructure deal.’ is less check-worthy than the claim ’Drinking salt water cures COVID.’.
+Not all identified claims are check-worthy.
+Given an input claim, the task is to identify the claims that include real-world assertions and that may need to be verified Hassan et al.
+(2015); Nakov et al.
+(2021).
+Check-worthy claim identification is a subjective task and it relies on factors like popularity of the claim, public interest to determine the claim’s veracity, etc.
+For example, the claim ’The President met the State Governor to discuss the infrastructure deal.’ is less check-worthy than the claim ’Drinking salt water cures COVID.’
 Claim Matching
-The identified check-worthy claims can be matched to the previously fact-checked claims.
-Given an input claim and a database of previously fact-checked claims, claim matching is used to determine if the input claim is previously fact-checked and exists in the database (Shaar et al.
-2020; Nakov et al.
-2021).
-This can help avoid the next steps in the pipeline and a veracity label can be predicted directly.
+The identified check-worthy claims can be matched to previously fact-checked claims.
+Given an input claim and a database of previously fact-checked claims, claim matching is used to determine if the input claim is previously fact-checked and exists in the database Shaar et al.
+(2020); Nakov et al.
+(2021).
+This can help predict the veracity label directly without additional steps.
 Document/Evidence Retrieval
-If the input claim does not exist in the database of fact-checked claims, the claim needs to be verified.
-In the document or evidence retrieval sub-task, all the relevant documents related to the input claim are extracted, either from an external database or through an internet search/information retrieval(Chen et al.
-2017).
-A threshold of how many documents are to be retrieved can be predetermined.
+If the input claim does not exist in the database of fact-checked claims, it needs to be verified.
+In the evidence retrieval sub-task, all relevant documents related to the input claim are extracted, either from an external database or through an internet search Chen et al.
+(2017).
+The threshold of the number of documents to be retrieved can be predetermined.
 Rationale/Sentence Selection
-All the information in the retrieved documents is not relevant to the claim.
-Hence in the rationale selection task, only the information or evidence most relevant to the claim are selected to be used to predict the veracity label.
+Not all information in the retrieved documents is relevant to the claim.
+Hence for the rationale selection task, only the information or evidence most relevant to the claim is selected to predict the veracity label.
 Veracity Label Prediction
-Once the rationale is selected, the claim along with the rationale and additional features if any, are given as input to a machine learning model.
-The task of the classifier is to predict a veracity label from among the following three labels; ’SUPPORTED’, ’REFUTED’, or ’NOT ENOUGH EVIDENCE’.
-In some datasets, the labels can also be ’TRUE’, or ’FALSE’, depending on the specific task.
+Once a rationale is selected, it is provided to a classifier along with the claim and potentially additionally features, to predict a veracity label; often ‘SUPPORTED’, ‘REFUTED’, vs ‘NOT ENOUGH EVIDENCE’.
+For some datasets, the labels can be ‘TRUE’ vs ‘FALSE’.
 Explanation/Justification Generation
 Recent works have focused on generating explanations for the veracity labels prediction.
-This specific task is focused on generating natural language justifications or explanations for the prediction considering the claim and evidence to generate these explanations.
-LLM Approaches
-In this section, we describe the recent advancements that enable LLMs to be robust in fact verification scenarios.
+This specific task is focused on generating natural language justifications or explanations for the prediction, considering the claim and evidence.
+4 LLM Approaches
 Figure 3 shows an example pipeline that encapsulates multiple component modules (i.e., Evidence Retrieval, Prompt Creation, Transfer Learning, and LLM Generation) for verifying claims.
-Different from traditional fact verification pipelines that select evidence set for verifying claims, LLM-based claim verification conditions generated text based on the concatenated input claim and retrieved evidence.
-The ability of large, pre-trained LMs augmented with retrieval enables them to perform well on knowledge-intensive tasks such as text generation.
-Evidence Retrieval Strategies
-RAG models, which have been developed to combat the issue of hallucination in LLMs in knowledge-intensive tasks, have shown success in the scope of fact verification (Lewis et al.
-2020; Izacard et al.
-2023; Gao et al.
-2023; Guan et al.
-2023).
-Early work such as in Lewis et al.
-(2020) developed a framework that incorporates a retriever to retrieve evidence from an external database such as Wikipedia for conditionally generating veracity labels in fact verification.
-Authors in Izacard et al.
-(2023) demonstrate that RAGs perform well on the FEVER shared task (Thorne et al.
-2018b)) in few-shot settings, showing approximately 5% improvement over large-scale LLMs such as Gopher (Rae et al.
-2021) with significantly fewer parameters (i.e., 11B compared to 280B).
-Other works consider the optimization of either document ranking (Glass et al.
-2022; Chen et al.
-2022c; Hofstätter et al.
-2023) or input claims (i.e., queries) (Hang, Yu, and Tan 2024) as a crucial step for improving evidence retrieval for veracity labeling.
-Authors in Hofstätter et al.
+Different from traditional fact verification pipelines that select evidence sets for verifying claims, LLM-based claim verification conditions generated text based on the concatenated input claim and retrieved evidence.
+Retrieval-augmented LLMs have been shown to perform well on knowledge-intensive tasks such as text generation.
+Figure 3: LLM-based claim verification pipeline.
+This involves creating a prompt from the retrieved evidence and the input claim as input to the LLM to generate a label, sentence evidence, and/or explanation of its response.
+4.1 Evidence Retrieval Strategies
+RAG models, which have been developed to address hallucination in LLMs for knowledge-intensive tasks, have shown success for fact verification Gao et al.
+(2023); Guan et al.
+(2023).
+Early work such as Lewis et al.
+(2020) develop a framework that retrieves evidence from an external database such as Wikipedia for conditionally generating veracity labels.
+Izacard et al., Izacard et al.
+(2023) demonstrate that RAGs perform well on the FEVER shared task Thorne et al.
+(2018b)) in few-shot settings, showing 5% improvement over large-scale LLMs such as Gopher Rae et al.
+(2021) with significantly fewer parameters (i.e., 11B compared to 280B).
+Other works consider the optimization of either document ranking Glass et al.
+(2022); Chen et al.
+(2022c) or input claims (i.e., queries) Hang et al.
+(2024) as a crucial step for improving evidence retrieval for veracity labeling.
+Hofstätter et al., Hofstätter et al.
 (2023) use an autoregressive re-ranker to get the most relevant passages from the retriever.
 These are then passed to the generation model to generate the veracity label.
-Despite this, RAG models can often fail when encountering long or complex input claims, causing the model to incorrectly generate veracity labels or evidence sentences.
-Recent works have addressed this issue by segmenting long claims into smaller sub-claims and performing multiple rounds of retrieval (Khattab, Potts, and Zaharia 2021; Shao et al.
-2023).
-Authors in Khattab, Potts, and Zaharia (2021) present a pipeline for multi-hop claim verification that uses an iterative retriever and neural methods for effective document retrieval and re-ranking.
+Despite this, RAG models often fail when encountering long or complex input claims, resulting in incorrectly generated veracity labels or evidence sentences.
+Recent works have addressed this by segmenting long claims into smaller sub-claims and performing multiple rounds of retrieval.
+Khattab et al., Khattab et al.
+(2021) presents a pipeline for multi-hop claim verification that uses an iterative retriever and neural methods for effective document retrieval and re-ranking.
+Shao et al.
 Shao et al.
 (2023) show that using a re-ranker to distill knowledge to a retriever helps close the semantic gaps between a query and document passage when verifying claims using iterative RAGs.
-Other works address the issue of complex claims by using fine-grained retrieval techniques based on claim decomposition (Chen et al.
-2023a; Zhang and Gao 2023; Pan et al.
-2023b).Chen et al.
-(2023a) generate sub-questions based on a claim, which a document retriever uses to retrieve relevant documents.
+Other works address the issue of complex claims using fine-grained retrieval techniques based on claim decomposition.
+Ches et al.
+Chen et al.
+(2023b) generate sub-questions based on a claim, which a document retriever uses to retrieve relevant documents.
 A fine-grained retriever retrieves top-k text spans as evidence based on a k-word window and BM25.
-Authors in Zhang and Gao (2023) decompose a claim into sub-claims and generate questions to verify the sub-claims.
+Zhang and Gao Zhang and Gao (2023) decompose a claim into sub-claims and generate questions to verify the sub-claims.
 External knowledge sources are used to retrieve relevant information to verify the sub-claims and generate a final veracity label.
+Pan et al.
 Pan et al.
@@
 The fact-verification subtask uses external knowledge to retrieve relevant evidence for a claim.
-Hang, Yu, and Tan (2024) retrieve evidence based on generated knowledge graphs of evidences.
-They generate a knowledge graph of user query or input claim and compare it to the database of knowledge graphs to retrieve the most relevant information for claim verification.
-Authors in Hu et al.
+Hang et al.
+Hang et al.
+(2024) retrieve evidence based on generated knowledge graphs of evidences.
+They generate a knowledge graph of the user query or input claim and compare it to the database of knowledge graphs to retrieve the most relevant information for claim verification.
+Hu et al.
+Hu et al.
 (2023) propose a latent variable model that allows the retrieval of the most relevant evidence sentences from a document while removing the irrelevant sentences.
-This approach reduces the noisy data during the verification process.
-Stochastic-RAG, an approach proposed by Zamani and Bendersky (2024) uses a stochastic sampling without replacement process for evidence retrieval and selection.
+This approach reduces noisy data during the verification process.
+Stochastic-RAG proposed by Zamani and Bendersky Zamani and Bendersky (2024) uses a stochastic sampling without replacement process for evidence retrieval and selection.
 This approach overcomes the ranking and selection of the evidence hence optimizing the RAG model.
-Authors in Zhang et al.
+Zhang et al.
+Zhang et al.
 (2023) optimize the evidence retrieval process by using feedback from the claim verifier.
@@
 Xu et al.
-(2024b) propose Search-in-the-Chain, an approach where an LLM generates a reasoning chain and based on the answer to each node in the chain, retrieval can be used to correct an answer or provide additional knowledge.
-The approach improves the generation accuracy of the LLM.
-Prompt Creation Strategies
-Text prompting has been shown to be an effective technique for improving the desired output of large-scale generative models.
-In the context of claim verification, several works investigate prompting strategies, using both manual and automated techniques, for building more robust claim verification systems Zhang and Gao (2023); Li et al.
-(2023c); Zeng and Gao (2023); Chen et al.
-(2023b).
-Authors in (Zhang and Gao 2023) develop a hierarchical prompting technique that enables LLMs to verify multiple sub-claims using a step-by-step approach.
-The work in Li et al.
+Xu et al.
+(2024b) propose Search-in-the-Chain, where an LLM generates a reasoning chain and based on the answer to each node in the chain, the retrieval can be used to correct an answer or provide additional knowledge, improving the generation accuracy of an LLM.
+4.2 Prompt Creation Strategies
+Text prompting has shown to be effective for improving the output of LLMs.
+In the context of claim verification, several works investigate both manual and automated prompting strategies to increase robustness.
+Zhang and Gao Zhang and Gao (2023) develop a hierarchical prompting technique to verify multiple sub-claims using a step-by-step approach.
+Li et al.
+Li et al.
 (2023c) demonstrates a self-sufficient claim verification through prompting instructions on multiple language models.
-ProToCo (Zeng and Gao 2023) demonstrates improved claim verification performance of LLMs by leveraging a consistency mechanism to construct variants of the original claim-evidence pair prompt based on three logical relations (i.e., confirmation, negation, uncertainty).
-Authors in Chen et al.
-(2023b) develop a unified retrieval framework that employs discrete, continuous, and hybrid prompt strategies for adjusting to various knowledge-intensive tasks such as claim verification.
-Other works such as the FactualityPrompts (Lee et al.
-2022) framework test the output generations of LLMs given an input prompt and use an external database such as Wikipedia to calculate factuality and quality measures compared to the ground truth.
-Other works aim to improve the LLM’s reasoning abilities by appending the claims with evidence during prompting in the context of claim verification and text generation (Parvez 2024; Dougrez-Lewis et al.
-2024).
-Transfer Learning Strategies
+ProToCo Zeng and Gao (2023) demonstrates improved claim verification performance by leveraging a consistency mechanism to construct variants of the original claim-evidence pair prompt based on three logical relations (i.e., confirmation, negation, uncertainty).
+Chen et al.
+Chen et al.
+(2023a) develop a unified retrieval framework that employs discrete, continuous, and hybrid prompt strategies for adjusting to various knowledge-intensive tasks such as claim verification.
+The FactualityPrompts Lee et al.
+(2022) framework tests the output generations of LLMs given an input prompt and uses an external database such as Wikipedia to calculate factuality and quality measures compared to the ground truth.
+Other works aim to improve an LLM’s reasoning abilities by appending the claims with evidence during prompting for claim verification and text generation Parvez (2024); Dougrez-Lewis et al.
+(2024).
+4.3 Transfer Learning Strategies
 Fine-Tuning.
-Although recent studies show the success of pre-trained LLMs on zero- or few-shot tasks, they often fail to verify real-world claims given their limited internal knowledge.
-The success of fine-tuning has motivated recent work on claim verification (Chen et al.
-2022b; Pan et al.
-2021; Zeng and Zubiaga 2024).
-The work in Chen et al.
-(2022b) leverages a language model to fine-tune over an external corpus for retrieving passage titles and evidence sentences using constrained beam search.
-The results showed improved performance against the traditional fact verification pipeline over the FEVER dataset.
-Other work demonstrates that using GPT models to generate synthetic training data improves the performance of LLMs on various tasks such as fact checking (Tang, Laban, and Durrett 2024) and claim matching (Choi and Ferrara 2024).
-Authors in Pan et al.
-(2021) develop a pipeline for creating a fact verification dataset and fine-tuning a language model by leveraging passages from Wikipedia to generate QA pairs related to claim veracity.
-The authors show that it can improve state-of-the-art language models in zero-shot settings.
-The work in (Zeng and Zubiaga 2024) shows that using unlabelled pairwise data to increase the alignment between claim-evidence pairs shows significant improvement of LLM performance in few-shot claim verification tasks.
-In addition, other recent works leverage techniques such as reinforcement learning to fine-tune models for improving the veracity of claims and supporting evidence Zhang and Gao (2024); Huang et al.
+Although recent studies show the success of pre-trained LLMs on zero- or few-shot tasks, they often fail at verifying real-world claims given their limited internal knowledge.
+The success of fine-tuning has motivated recent work on claim verification.
+Chen et al.
+Chen et al.
+(2022b) fine-tuned an LM on an external corpus for retrieving passage titles and evidence sentences using constrained beam search, finding improved performance on the FEVER dataset.
+Other work demonstrates that using GPT models to generate synthetic training data improves the performance of LLMs on fact checking Tang et al.
+(2024) and claim matching Choi and Ferrara (2024).
+Pan et al.
+Pan et al.
+(2021) develop a pipeline for creating a fact verification dataset and fine-tuning a language model by leveraging passages from Wikipedia to generate QA pairs related to claim veracity, showing improved zero-shot performance.
+Zeng and Zubiaga Zeng and Zubiaga (2024) show that unlabelled pairwise data can increase the alignment between claim-evidence pairs, resulting in significant improvement on few-shot claim verification.
+Other recent works leverage reinforcement learning to fine-tune models for improving the veracity of claims and supporting evidence Zhang and Gao (2024); Huang et al.
 (2024).
-A document-level and question-level retrieval policy is proposed by Zhang and Gao (2024), where the top-k and top-1 documents for the document and question-level policy respectively are used as input to a scoring function for label prediction during training.
+A document-level and question-level retrieval policy is proposed by Zhang and Gao Zhang and Gao (2024), where the top-k and top-1 documents for the document and question-level policy respectively are used as input to a scoring function for label prediction during training.
 This approach outperforms retrieval and prompting approaches.
 Chiang et al.
-(2024) fine-tune LLMs for their multi-stage fact verification approach.
+Chiang et al.
+(2024) fine-tune LLMs for multi-stage fact verification.
 They fine-tune a model to generate answers based on claim-evidence pairs and a set of questions, whereas another model is fine-tuned to verify the claim based on the claim-evidence and question-answer pairs.
-Authors in Zhu et al.
-(2023) fine-tune a generation model to generate counterfactuals to train the fact verification model’s performance on out-of-domain claims.
+Zhu et al.
+Zhu et al.
+(2023) fine-tune a generation model to generate counterfactuals for out-of-domain classification.
 In-Context Learning.
-The recent success in the performance of pre-trained LLMs in zero- and few-shot settings is largely attributed to its in-context learning (ICL) abilities (Kojima et al.
-2022; Brown et al.
-2020).
-In the scope of claim verification, popular ICL techniques with strong zero- and few-shot performance include chain-of-thought (CoT) reasoning (Wei et al.
-2022).
-The work in Zhao et al.
-(2024) develops a multi-stage verification pipeline for claim verification based on claim decomposition and self-reflection.
+The recent success of pre-trained LLMs in zero- and few-shot settings is largely attributed to its in-context learning (ICL) abilities Kojima et al.
+(2022); Brown et al.
+(2020).
+For claim verification, popular ICL techniques include chain-of-thought (CoT) reasoning Wei et al.
+(2022).
+Zhao et al.
+Zhao et al.
+(2024) develop a multi-stage verification pipeline based on claim decomposition and self-reflection.
 An LLM-based verifier module is created using instruction prompting to generate a reasoning analysis among all sub-claims created by the decomposer module.
-The results suggest that using zero-shot prompting techniques provides better performance in multi-hop claim verification domains such as HOVER and FEVEROUS compared to few-shot prompting and fine-tuning methods.
-Authors in Kanaani (2024) enable LLMs to generate reasons over retrieved evidence in claim verification using few-shot ICL and the STaR CoT technique inspired by the work of Zelikman et al.
+The suggests that zero-shot prompting techniques result in better multi-hop performance on the HOVER and FEVEROUS datasets compared to few-shot prompting and fine-tuning.
+Kanaani Kanaani (2024) enable LLMs to generate reasons over retrieved evidence in claim verification using few-shot ICL and the STaR CoT technique inspired by Zelikman et al.
+Zelikman et al.
 (2022).
-Similar work has leveraged CoT techniques for the purposes of effectively verifying complex claims using reasoning steps (Yao et al.
-2023; Ni et al.
-2024).
-On the other hand, HiSS (Zhang and Gao 2023) demonstrates that using prompting techniques for few-shot learning and claim decomposition can substantially improve the performance of CoT models for complex news claim verification.
+Similar work has leveraged CoT for effectively verifying complex claims using reasoning steps Yao et al.
+(2023); Ni et al.
+(2024).
+Conversely, HiSS Zhang and Gao (2023) demonstrates that prompting for few-shot learning and claim decomposition can substantially improve the performance of CoT models for complex news claim verification.
+Li et al.
 Li et al.
@@
 Their approach outperforms or achieves comparable task performance in a zero-shot setting on claim verification datasets.
-LLM Generation Strategies
+4.4 LLM Generation Strategies
 Label and Evidence Generation.
-While the majority of claim verification systems predict veracity labels based on the concatenation of the input claim and the set of evidence sentences (Pradeep et al.
-2021), recent work has proposed alternate strategies for determining veracity labels as well as selecting/generating evidence pieces.
-The work in Cao et al.
-(2024) develops, SERIf, a claim verification pipeline that features an inference module that predicts the veracity label of scientific news articles based on a two-step summarization (i.e., ‘Extractive - Abstractive‘) and evidence retrieval technique.
-Each summary-evidence pair is fed into the LLM and it produces a binary label, indicating whether the news article is reliable (supported) or unreliable (refuted).
-Authors in Wadden et al.
-(2022b) leverage the Longformer transformer model (Beltagy, Peters, and Cohan 2020) that uses a shared encoding over the claim and document abstracts for rationale identification and claim label prediction.
-The work in Li et al.
-(2024a) develops an algorithm that selects the minimal evidence group (MEG) within a set of retrieved candidate documents.
-The algorithm aims to minimize the redundancy while also selecting the most relevant piece of evidence to prompt the language model.
-Authors in Chen et al.
-(2022b) create an LLM claim verification framework in which the authors use the BART model to encode all candidate sentences from the most relevant retrieved documents.
-The BART decoder serves as an evidence decoder to predict the $g$-th evidence sentence via generation conditioned on the top $k$ retrieved documents and the input claim.
+While the majority of claim verification systems predict veracity based on the concatenation of the input claim and evidence sentences Pradeep et al.
+(2021), recent work has proposed alternate strategies for determining veracity labels and selecting/generating evidence pieces.
+Cao et al.
+Cao et al.
+(2024) develop SERIf, a claim verification pipeline that features an inference module to predict the veracity label of scientific news articles based on a two-step summarization (i.e., ‘Extractive - Abstractive‘) and evidence retrieval technique.
+Each summary-evidence pair is fed into the LLM and produces a binary label, indicating whether the news article is reliable (supported) or unreliable (refuted).
+Wadden et al.
+Wadden et al.
+(2022b) leverage the Longformer model Beltagy et al.
+(2020) that uses a shared encoding over the claim and document abstracts for rationale identification and claim label prediction.
+Li et al.
+Li et al.
+(2024a) propose to select the minimal evidence group within a set of retrieved candidate documents.
+This aims to minimize redundancy while selecting the most relevant evidence to prompt the language model.
+Chen et al.
+Chen et al.
+(2022b) use BART to encode all candidate sentences from the most relevant retrieved documents.
+In this, BART serves as an evidence decoder to predict the $g$-th evidence sentence via generation conditioned on the top $k$ retrieved documents and the input claim.
 Lee et al.
-(2022) developed a variant of nucleus sampling called, factual-nucleus sampling, in which the top-$p$ sampling pool is selected as a set of sub-words whose cumulative probability exceeds $p$.
-The authors show that factual-nucleus sampling can improve evidence and label generation without claim verification datasets such as FEVER.
-Kao and Yen (2024) propose a multi-stage approach, where the evidence sentences are retrieved from articles related to a claim, and arguments are generated by aggregating and reconstructing the evidence.
+Lee et al.
+(2022) develop a variant of nucleus sampling called factual-nucleus sampling, in which the top-$p$ sampling pool is selected as a set of sub-words whose cumulative probability exceeds $p$, resulting in improved evidence and label generation without claim verification datasets such as FEVER.
+Kao and Yen Kao and Yen (2024) propose a multi-stage approach, where the evidence sentences are retrieved from articles related to a claim, and arguments are generated by aggregating and reconstructing the evidence.
 The arguments are refined and passed to a LLM to generate a verification label.
-Other approaches leverage LLMs reasoning capabilities to generate veracity labels and factual evidence (Cheng, Tan, and Lu 2024; Jafari and Allan 2024; Li et al.
-2024b; Fang et al.
-2024; Pan et al.
-2023a).
-Dataset # Instances Source # Labels Reference
-AVeriTeC 4,568 Google FactCheck Claim Search API 4 Schlichtkrull, Guo, and Vlachos (2024)
-BINGCHECK 3,840 Microsoft Copilot 4 Li et al. (2023c)
-Check-COVID 1,504 News websites, CORD-19 3 Wang et al. (2023a)
-CLAIMDECOMP 1,000 PolitiFact 6 Chen et al. (2022a)
-CLIMATE-FEVER 1,535 Google Search, Wikipedia 3 Diggelmann et al. (2020)
-CoVERT 212 Manually created 3 Liu et al. (2024)
-COVID-Fact 4,086 Reddit 2 Saakyan, Chakrabarty, and Muresan (2021)
-EX-FEVER 60,000 Wikipedia 3 Ma et al. (2024)
-ExpertQA 2177 LLM, Google Search 5 Malaviya et al. (2024)
-FACTIFY-5WQA 391,041 Public Fact-Verification Datasets 3 Rani et al. (2023)
-FactCheck-Bench 678 LLM 4 Wang et al. (2023b)
-FactCheckQA 20,871 PolitiFact 3 Bashlovkina et al. (2023)
-FACTUALITYPROMPTS 16,000 Wikipedia, LLM 2 Lee et al. (2022)
-FAVIQ 188,000 Wikipedia, Existing Datasets 2 Park et al. (2022)
-FEVER 185,445 Wikipedia 3 Thorne et al. (2018a)
-FEVEROUS 87,026 Wikipedia 3 Aly et al. (2021)
-FoolMeTwice 12,968 Wikipedia 2 Eisenschlos et al. (2021)
-HealthVer 14,330 Bing Web Search API 3 Sarrouti et al. (2021)
-HOVER 26,171 Wikipedia 2 Jiang et al. (2020)
-LIAR 12,500 PolitiFact 6 Wang (2017)
-MultiFC 34,918 Fact-checking Websites 2-40 Augenstein et al. (2019)
-PolitiFact 21,152 PolitiFact 6 Misra (2022)
-QuanTemp 15,514 Google Fact Check Tool API 3 Venktesh et al. (2024)
-RAWFC 2,012 Snopes 6 Yang et al. (2022)
-SciFact 1,400 S2ORC 2 Wadden et al. (2020)
-SciFact-Open 279 S2ORC 3 Wadden et al. (2022a)
-SciNews 2,400 Wiki Fake News, LLM, Existing Datasets 2 Cao et al. (2024)
-SCITAB 1,225 SciGen Dataset (arXiv) 3 Lu et al. (2023)
-VitaminC 400,000 Wikipedia 3 Schuster, Fisch, and Barzilay (2021)
-WiCE 1,967 Wikipedia 3 Kamoi et al. (2023)
-XClaimCheck 16,177 PolitiFact 5 Kao and Yen (2024)
-Table 1: Available datasets along with the number of instances, data sources, and references.
+Other approaches leverage LLMs reasoning capabilities to generate veracity labels and factual evidence Cheng et al.
+(2024); Jafari and Allan (2024); Li et al.
+(2024b); Fang et al.
+(2024); Pan et al.
+(2023a).
 Explainable Generation.
-Recent studies investigate explainable approaches to improve LLM-based claim verification systems (Wang and Shu 2023; Dammu et al.
-2024; Si et al.
-2023).
-Authors in Wang and Shu (2023) present an explainable claim verification framework named FOLK, that leverages the explanation capabilities of LLMs when verifying a claim and justifies the prediction through a summary of its decision process.
-The work in Dammu et al.
-(2024) proposes a knowledge graph (KG)-based approach for text verification and evidence attribution.
-An objective function is used to fine-tune LLMs on evidence attribution based on the input text and retrieved triplets from KG, inducing explanations on claim predictions.
-While explainable techniques can aid humans in fact-checking, LLMs are prone to incorrect explanations due to hallucinations, causing them to be unreliable in certain claim verification scenarios (Si et al.
-2023).
+Recent studies investigate explainable approaches to improve LLM-based claim verification.
+Wang and Shu Wang and Shu (2023) present the FOLK framework, that leverages the explanation capabilities of LLMs when verifying claims and justifies the prediction through a summary of its decision process.
+Dammu et al.
+Dammu et al.
+(2024) proposes a knowledge graph (KG)-based approach for text verification and evidence attribution, where the LLM is fine-tuned on evidence attribution based on the input text and retrieved triplets from KG, inducing explanations for claim predictions.
+While explanation techniques can help humans verify facts, LLMs can produce incorrect explanations due to hallucinations, making them unreliable in certain claim verification scenarios Si et al.
+(2023).
 Ma et al.
-(2024) prompt an LLM in a few-shot setting to generate a concise summary of the gathered evidence documents and the input claim.
-This summary serves as an explanation for the verified claim.
-Other works in claim verification leverage reasoning techniques such as chain-of-thought (CoT) for enabling the LLM to be interpretable in its decision-making process when verifying claims (Yao et al.
-2023; Pan et al.
-2023a; Zhao et al.
-2024; Kanaani 2024; Ni et al.
-2024; Quelle and Bovet 2024; Fang et al.
-2024).
-The authors in Pan et al.
+Ma et al.
+(2024) prompt an LLM in a few-shot setting to generate a concise summary of the evidence documents and input claim, serving as an explanation for the verified claim.
+Other works leverage reasoning techniques such as chain-of-thought (CoT) to enable the LLM to be interpretable in its decision-making process when verifying claims Yao et al.
+(2023); Pan et al.
+(2023a); Zhao et al.
+(2024); Kanaani (2024); Ni et al.
+(2024); Quelle and Bovet (2024); Fang et al.
+(2024).
+Pan et al.
+Pan et al.
 (2023a) and Fang et al.
-(2024) leverage the reasoning ability of the LLMs to generate explanations by using question-guided reasoning and minimizing the inherent model biases respectively.
-Evaluation and Benchmarking
-Metrics
-The F1 score is the most commonly used metric to measure the performance of automatic claim verification systems.
-Other metrics like Precision, Recall, and Accuracy are also used to evaluate the system performance.
-Katranidis and Barany (2024) used the error rate between the human and automated fact-verification system to measure the verification accuracy.
+Fang et al.
+(2024) leverage an LLM’s reasoning ability to generate explanations by using question-guided reasoning and minimizing the inherent model biases.
+5 Evaluation and Benchmarking
+5.1 Metrics
+The F1 score is the most commonly used metric to measure the performance of automatic claim verification systems, and to a lesser degree Precision, Recall, and Accuracy.
+Katranidis and Barany Katranidis and Barany (2024) used the error rate between the human and automated fact-verification system to measure verification accuracy.
 However, these metrics consider a single pipeline component to evaluate the system’s overall performance.
 Hence, Thorne et al.
-(2018a) introduced the FEVER score, a metric that uses both the verification accuracy as well as the evidence retrieval accuracy to compute overall system performance.
+Thorne et al.
+(2018a) introduce the FEVER score, a metric that incorporates both verification accuracy and evidence retrieval accuracy to compute overall system performance.
 While these metrics are valuable for assessing the performance of the classification tasks, they are inadequate for evaluating the performance of the non-classification components of the pipeline.
-Hence, metrics like Recall@k are used to measure the performance of the retrieval task (Pan et al.
-2023b; Pradeep et al.
-2021), while BLEU and METEOR are used to evaluate the quality of explanations or generated questions and answers.
-Schlichtkrull, Guo, and Vlachos (2024) propose a new evaluation metric AVeriTeC score that uses METEOR and accuracy metrics, for question-answer-based veracity label prediction systems.
+Hence, metrics like Recall@k are used to measure the performance of the retrieval task Pan et al.
+(2023b); Pradeep et al.
+(2021), while BLEU and METEOR are used to evaluate the quality of explanations or generated questions and answers.
+Schlichtkrull et al.
+Schlichtkrull et al.
+(2024) propose the new evaluation metric AVeriTeC score that uses METEOR and accuracy for question-answer-based veracity prediction systems.
 Other metrics like Mean Absolute Error (MAE), Expected Calibration Error (ECE), Area Under ROC Curve (AUC-ROC), and Pearson’s Correlation are also used.
-Most of these metrics are inadequate to evaluate the factual accuracy of the LLM-generated text.
-Several metrics like FactScore (Min et al.
-2023), SAFE (Wei et al.
-2024), and VERISCORE (Song, Kim, and Iyyer 2024) are introduced to evaluate the factual accuracy of the LLM-generated text.
-While these metrics evaluate the accuracy, other metrics and frameworks to evaluate the factual errors in generated text (Lee et al.
-2022; Chern et al.
-2023) as well as their alignment (Zha et al.
-2023) and entailment (Lee et al.
-2022) considering the factuality have been proposed.
-Datasets
+Most of these metrics are inadequate for evaluating the factual accuracy of LLM-generated text, for which FactScore Min et al.
+(2023), SAFE Wei et al.
+(2024), and VERISCORE Song et al.
+(2024) have been proposed instead.
+Other metrics and frameworks evaluate factual errors in generated text Lee et al.
+(2022); Chern et al.
+(2023) and their alignment Zha et al.
+(2023) and entailment Lee et al.
+(2022) considering factuality.
+5.2 Datasets
 The fundamental resource for training and evaluating claim verification systems is datasets containing annotated texts.
-As most research in this area deals with English data, we collect information about all publicly available English datasets used in the papers discussed in this survey and present them in Table 1.
-Several general-domain datasets have been released over the years.
-Different online data sources and websites are used to extract facts and claims including Wikipedia (Thorne et al.
-2018a; Jiang et al.
-2020; Diggelmann et al.
-2020; Eisenschlos et al.
-2021; Schuster, Fisch, and Barzilay 2021; Kamoi et al.
-2023), due to the extensive amount of information available spanning various topics and domains, and online fact-checking websites like PolitiFact (Wang 2017; Augenstein et al.
-2019; Kao and Yen 2024).
+As most research in this area deals with English data, we collect information about all publicly available English datasets used in the papers discussed in this survey [^4] https://github.com/LanguageTechnologyLab/Claim-Verification-Papers.git.
+General-domain datasets have been created from online data sources and websites including Wikipedia Thorne et al.
+(2018a); Jiang et al.
+(2020); Diggelmann et al.
+(2020); Eisenschlos et al.
+(2021); Schuster et al.
+(2021); Kamoi et al.
+(2023), due to the extensive amount of information available spanning various topics and domains, and online fact-checking websites like PolitiFact Wang (2017); Augenstein et al.
+(2019); Kao and Yen (2024).
 Factually incorrect claims are shared through social media channels like X, Reddit, etc.
-Saakyan, Chakrabarty, and Muresan (2021) introduced the COVID-Fact dataset consisting of claims extracted from Reddit posts.
-Several datasets for scientific fact verification have also been introduced over the years (Wadden et al.
-2020, 2022a; Lu et al.
-2023).
+Saakyan et al.
+Saakyan et al.
+(2021) introduce the COVID-Fact dataset consisting of claims extracted from Reddit posts.
+Several datasets for scientific fact verification have also been introduced Wadden et al.
+(2020, 2022a); Lu et al.
+(2023).
 Most fact-verification datasets rely on unstructured textual evidence.
-Hence, a few datasets with structured data sources as evidence have been introduced (Aly et al.
-2021; Lu et al.
-2023).
-While a large number of datasets are focused on only veracity labels, some datasets were deloped to address explainability as well (Chen et al.
-2022a; Yang et al.
-2022; Ma et al.
-2024; Rani et al.
-2023; Schlichtkrull, Guo, and Vlachos 2024).
-Most of the claim-verification datasets include claims extracted from available information sources.
-LLMs are widely used to generate content, given their superior text-generation capabilities.
-These models often generate factually incorrect information.
+Hence, a few datasets with structured evidence sources have been introduced Aly et al.
+(2021); Lu et al.
+(2023).
+While many datasets focus only on veracity labels, some were developed to address explainability too Chen et al.
+(2022d); Yang et al.
+(2022); Ma et al.
+(2024); Rani et al.
+(2023); Schlichtkrull et al.
+(2024).
+Most claim-verification datasets include claims extracted from available information sources.
+LLMs are widely used to generate content, even though they can generate factually incorrect information.
 Efforts to identify such non-factual text have been undertaken.
-However, existing claim-verification datasets can be inadequate for this task due to the linguistic variations between the human- and LLM-generated text.
-To overcome this issue, LLM-generated claim verification datasets have been introduced (Li et al.
-2023c; Cao et al.
-2024).
-While these datasets are used to evaluate the automatic claim verification systems, there is a need to evaluate the factual accuracy of the LLM-generated text.
-Hence, a few datasets to evaluate the LLM generation’s factual accuracy have been introduced (Lee et al.
-2022; Wang et al.
-2023b; Malaviya et al.
-2024).
-Shared Tasks
-Shared tasks are competitions where participating teams develop systems to solve a task using a common benchmark dataset.
-There have been multiple shared tasks on the claim and fact-checking organized over the years like the Fact Extraction and Verification (FEVER) (Thorne et al.
-2018b), TabFact [^4] https://competitions.codalab.org/competitions/21611, CLEF 2020 CheckThat!
-(Barrón-Cedeno et al.
-2020), SCIVER (Wadden and Lo 2021), SEM-TAB-FACT (Wang et al.
-2021), FACTIFY-5WQA [^5] https://defactify.com/factify3.html, and more recently AVeriTeC [^6] https://fever.ai/task.html.
-While there have been various techniques like question-answer generation as a precursor task to label prediction as in the AVeriTeC shared task, all these shared tasks are centered on predicting a veracity label for a claim.
-Given the LLM’s tendency to hallucinate generating plausible yet factually inaccurate text, there is a need to organize shared tasks to evaluate the factual accuracy of LLM-generated content.
-Open Challenges
+However, claim-verification datasets consisting of human-written claims can be inadequate for this task due to the linguistic discrepancy between human- and LLM-generated text.
+As such, LLM-generated claim verification datasets have been introduced Li et al.
+(2023c); Cao et al.
+(2024).
+While these are used to evaluate automatic claim verification systems, there is a need to evaluate the factual accuracy of LLM-generated text, leading to the introduction of evaluation datasets Lee et al.
+(2022); Wang et al.
+(2024); Malaviya et al.
+(2024).
+5.3 Shared Tasks
+Shared tasks are competitions where teams develop systems for tasks using a common benchmark dataset.
+Multiple shared tasks on claim and fact-checking have been organized including the Fact Extraction and Verification (FEVER) challenge Thorne et al.
+(2018b), TabFact [^5] https://competitions.codalab.org/competitions/21611, CLEF 2020 CheckThat!
+Barrón-Cedeno et al.
+(2020), SCIVER Wadden and Lo (2021), SEM-TAB-FACT Wang et al.
+(2021), FACTIFY-5WQA [^6] https://defactify.com/factify3.html, and more recently AVeriTeC [^7] https://fever.ai/task.html.
+While there have been various techniques like question-answer generation as a precursor to label prediction as in AVeriTeC, all these shared tasks are centered on predicting veracity labels.
+Given an LLM’s tendency to generate plausible yet factually inaccurate text, there is a need to organize shared tasks to evaluate the factual accuracy of LLM-generated content.
+6 Open Challenges
 Handling Irrelevant Context
 Retrieved evidence may be irrelevant, which is a challenge for LLMs, as they may not be trained to ignore such evidence.
-The lack of robustness to the noise can cause the LLM to produce misinformation and incorrect verification.
-Recent research on open-domain question answering shows that external knowledge relevant to the task can aid the model performance, however, irrelevant context can also lead the model to make inaccurate predictions (Petroni et al.
-2020; Shi et al.
-2023; Li et al.
-2023a; Yu et al.
-2023).
-For the fact-verification task, recent works have proposed techniques to identify the most relevant context while aiding the veracity label prediction and thus improving the overall system performance (Wang et al.
-2023c; Yoran et al.
-2024; Xia et al.
-2024).
-However, more work on the topic is required to verify the effectiveness of this design approach across multiple fact verification domains.
+The lack of robustness to noise can lead an LLM to produce misinformation and incorrect verification.
+Recent research on open-domain question answering shows that external knowledge relevant to the task can improve model performance, however, irrelevant context can also lead to inaccurate predictions Petroni et al.
+(2020); Shi et al.
+(2023); Li et al.
+(2023a); Yu et al.
+(2023).
+For the fact-verification, recent work has proposed techniques to identify the most relevant context for improving veracity label prediction, thus improving the overall system performance Wang et al.
+(2023); Yoran et al.
+(2024); Xia et al.
+(2024).
+However, more work is required to verify the effectiveness across domains.
 Handling Knowledge Conflicts
-Fact-verification approaches reliance on retrieved evidence can cause knowledge conflicts in LLM-based approaches where retrieved evidence as the internal parameters of the pre-trained LLM may conflict with the external knowledge.
-This causes the LLM to ignore the retrieved evidence and produces hallucinations.
-Xu et al.
-(2024a) provides an in-depth analysis of this scenario.
-Works by (Li et al.
-2023a; Neeman et al.
-2023; Mallen et al.
-2023; Longpre et al.
-2021; Chen, Zhang, and Choi 2022) introduce approaches to avoid knowledge conflicts and mitigate hallucinations for question-answering.
+The reliance of fact-verification approaches on retrieved evidence can cause knowledge conflicts in LLMs, where retrieved external evidence may conflict with the internal parameters of the pre-trained LLM.
+This causes the LLM to ignore the retrieved evidence and produces hallucinations Xu et al.
+(2024a).
+Approaches for avoiding knowledge conflicts have been introduced for question answering  Li et al.
+(2023a); Neeman et al.
+(2023); Mallen et al.
+(2023); Longpre et al.
+(2021); Chen et al.
+(2022a); Marjanović et al.
+(2024).
 Expanding this work to fact-verification, especially to the approaches that use LLMs, is vital for an effective verification process.
@@
 Most automated claim verification approaches rely on English datasets.
-Furthermore, there are limited multilingual fact-verification datasets (Gupta and Srikumar 2021; Kazemi et al.
-2022; Pikuliak et al.
-2023; Singh et al.
-2023).
-This hinders the development of approaches for multilingual fact-verification, which achieve the best performance when trained on language-specific datasets (Panchendrarajan and Zubiaga 2024).
-Conclusion
+Furthermore, there are limited multilingual fact-verification datasets Gupta and Srikumar (2021); Kazemi et al.
+(2022); Pikuliak et al.
+(2023); Singh et al.
+(2023).
+This hinders the development of approaches for multilingual fact-verification, which achieve the best performance when trained on language-specific datasets Panchendrarajan and Zubiaga (2024).
+7 Conclusion
 We presented a survey on LLM approaches to claim verification.
-To the best of our knowledge, this is the first claim verification survey to focus exclusively on LLM approaches thus filling an important gap in the literature.
-We have described each of the sub-tasks of the typical claim verification pipeline and discussed various LLM-based approaches used in this task.
+To the best of our knowledge, this is the first claim verification survey to focus exclusively on LLM approaches, thus filling an important gap in the literature.
+We have described the traditional claim verification pipeline including its component tasks and discussed various LLM-based approaches used in this task.
 Finally, we have also described publicly available English datasets providing important information to new and seasoned researchers on this topic.
-Advances in LLM development will likely continue improving the quality of claim verification systems.
-We hope this survey motivates future research on this topic taking advantage of recently-proposed LLMs, RAG methods, etc.
-Claim verification is a vibrant research topic and we see multiple open research directions as described in the next sub-section.
+Advances in LLM development will likely continue to improve the quality of claim verification systems.
+We hope this survey motivates future research on this topic taking advantage of recently-proposed LLMs, RAG methods, and others.
+Claim verification is a vibrant research topic and we envisage multiple open research directions such as handling irrelevant context, knowledge conflicts and multilingualism.